OpenStudy (anonymous):

What is the weight of a mass of 0.001kg on the surface of a neutron star when the gravitational field strength is 2.3*10^11 N kg?

OpenStudy (anonymous):

@DarrelFletcher, Please make sure that you are careful while posting the right question. gravitational field strength--> g = F /m =N kg^(-1) Your question says 2.3*10^11 N kg . The units are wrong :( Answer: g = 2.3 * 10^11 N Kg^(-1) F= mg Hope you can solve it from here

OpenStudy (anonymous):

The answer is 230000000 Newtons how do i scale that down to the appropriate answer?

OpenStudy (anonymous):

2.3*10^8 N

OpenStudy (anonymous):

ok thanks